OpenStudy (anonymous):

If f(x) = 4x + 1 and g(x) = x2 – 2x + 1, what is f(g(x)) ?

OpenStudy (lgbasallote):

substitute g(x) to your x in f(x) so it's \[4(x^2 - 2x +1) +1\]

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow f(g(x)) = f(x^2 - 2x + 1) = 4(x^2 - 2x + 1) + 1 }\)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

but these are the choices they give 4x2 – 8x + 5 4x2 + 2x + 2 x2 + 2x + 2 x2 – 2x + 2 which one is correct

OpenStudy (lgbasallote):

simplify the expression i gave to find out

OpenStudy (lgbasallote):

just distribute 4 into each term of \(x^2 - 2x +1\)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Find a third–degree polynomial function such that f(0) = –12 and whose zeros are 1, 2, and 3. Using complete sentences, explain how you found it.